the base of a triangle is b. the height is 3 more than half the base.
a. find the area if the base is 10.

Answer by checkley71(8403)

AREA=1/2BH
AREA=1/2*B(B/2+3)
AREA=1/2*10(10/2+3)
AREA=5*10(5+3)
AREA=5*10*8
AREA=400 ANSWER.